Unsure of my answer for c.
Multiple-choice questions each have four possible answers (a, b, c, d), one of which is correct. Assume that you guess the answers to three such questions. rect answer and W denotes a wrong answer Use the multiplication rule to find P(WCC), where C denotes a cor P/Wcc)- [046875 Type an exact answer) b. Beginning with WCC, make a complete list of the different possible arrangements of two correct answers and one wrong answer, then find the probability for each entry in the list P(WCC)-see above P(ccW) 046875 P(CWC) 046875 Type exact answers.) c. Based on the preceding results, what is the probability of getting exactly two correct answers when three guesses are made? 1.0001031 (Type an exact answer.)
Answer #1

c.

We will add all probabilities

P(exactly 2 out of 3 correct) = P(WCC)+P(CWC)+P(CCW)

= 0.046875+0.046875+0.046875 = 0.140625